2.  Suppose that a hypothetical “consumer market basket” consists only

of goods B and C, in the quantities:  B = 10 and C = 5.  

Use 2018 as a base year (i.e., 2018 = 100).

                                                                   Year 2017      Year 2018     Year 2019

Quantity of Good A                                            3                      4                     5

Price of Good A                                                 $9                  $10                $11

Quantity of Good B                                          10                    10                   10

Price of Good B                                                 $2                    $4                   $6

Quantity of Good C                                            2                      4                      6

Price of Good C                                                 $5                    $6                    $7

a.  What is the total money value spent on the consumer market basket in 2018 and 2019?

b.  Calculate the CPI for 2018, and 2019.

c.  What is the inflation rate for 2017 - 2018?  

d.  What is the inflation rate for 2018 - 2019?
Answer #1

Given that consumer market basket consist only B & C.
Also specified that quantity of B = 10 & C = 4

a) Total money value of a basket is the revenue that can be earned from the basket.

  • Total money value of commodity basket in 2018 of B = Price of B * Quantity of B
      = $4 * 10 = $40
  • Total money value of commodity basket in 2019 of B = Price of B * Quantity of B
    = $6 * 10 = $60
  • Total money value of commodity basket in 2018 of C = Price of B * Quantity of B
    = $6 * 4 = $24
    since in 2019 , quantity of C is not 4 hence it is not consumer basket and hence money value is not calculated.
  • Total money value of 2018 = $40 + $24 = $64
    Total money value of 2019 = $60

b) Given that base year is 2018
Rewritting the prices for each good as follows:

GOOD 2017 2018BASE 2019
A $9 $10 $11
B $2 $4 $6
C $5 $6 $7
TOTAL : $16 $20 $24

CPI 2017 : (TOTAL PRICE IN 2017 / TOTAL PRICE OF BASE YEAR)* 100
CPI = ($16 / $20) * 100
= 0.8 * 100 = 80

CPI 2018 : (TOTAL PRICE IN 2018 / TOTAL PRICE OF BASE YEAR)* 100
CPI = ($20 / $20) * 100
= 1 * 100 = 100

CPI 2019 : (TOTAL PRICE IN 2019 / TOTAL PRICE OF BASE YEAR)* 100
CPI = ($24 / $20) * 100
= 1.2 * 100 = 120

CPI FOR 2018 = 100
CPI FOR 2019 = 120

C) Inflation rate between 2017 - 2018 :
Inflation rate = [(CPI 2018 - CPI 2017)/CPI 2017] * 100
= [ (100 - 80)/80] * 100
= [ 20/80] * 100
= 0.25*100 = 25%
Inflation = 25%

d) Inflation rate between 2018 - 2019 :
Inflation rate = [(CPI 2019 - CPI 20178)/CPI 2018] * 100
= [ (120 - 100)/100] * 100
= [ 20/100] * 100
= 20%
Inflation = 20%